OpenStudy (anonymous):

please help :{ Solve p =w/t for t a. t = p - w c. t=w/p b. t = 2 + p d. t=p/w

OpenStudy (jack1):

... which part do you need help with...?

OpenStudy (jack1):

@leahhbe ...?

OpenStudy (anonymous):

like i just dont understand the problem i just got on this lesson @Jack1

OpenStudy (anonymous):

t=w/p

OpenStudy (jack1):

"t=w/p" yep from p = w/t multiply both sides of the equals sign by t p x t = w/t x t p t = w x 1 (as t/t = 1) so p t = w then divide both sides by p (p t) / p = w / p t x 1 = w / p (as p/p = 1) so t = w / p